when john was marys age she was 7. when shes his age he will be 73. What is John's age?
Answer by ankor@dixie-net.com(22740)

when john was marys age she was 7.
when shes his age he will be 73.
What is John's age?
:
let j = John's present age
let d = the difference in their ages
:
"when John was Mary's age she was 7.
j = d + 7
or
j - d = 7
"when she's his age he will be 73."
j + d = 73
Use elimination with these two equations
j - d = 7
j + d = 73
2j = 80
j = 80/2
j = 40 yrs is John's age
: